Both epsilon 0 and Xe are dimensionless.

(a) True

(b) False


1
Expert's answer
2021-07-21T08:34:01-0400

(1) "\epsilon_0"

Dimensions

"F=\frac{Kq_1q_2}{r^2}=\frac{q_1q_2}{4\pi \epsilon_0 r^2}"

"\epsilon_0=\frac{4\pi Fr^2}{q_1q_2}\rightarrow(1)"

Find dimensions

Put value equation (1)

"\epsilon_0=\frac{[M][LT^{-2}][L^2]}{[AT]^2}"

"\epsilon_0=[ML^3T^{-4}A^{-1}]"

"\epsilon_0" Exist dimensions

(2)

"\chi" dimensions

"\chi=\frac{M}{H}\rightarrow(2)"

equation (2) put value

H and M dimensions put find magnetic susceptibility dimensions

"\chi=[M^1L^1T^{-2}A^{-2}]"

"\chi" exist dimensions

Both eqution exist dimensions

Statement are false

Option (b) is correct option